OpenStudy (conqueror):

Can someone please tell me a little about Linear Functions and Linear Models?

OpenStudy (conqueror):

@SolomonZelman

OpenStudy (conqueror):

@iGreen

OpenStudy (igreen):

Linear Functions have straight lines.

OpenStudy (igreen):

Example 

OpenStudy (igreen):



OpenStudy (michele_laino):

here is something on linear functions: linear function or homorphysms, are function acting between vector spaces. Suppose that f(x) is a linear function, acting between the subsequent vector spaces: V, and W now, by definition, we have: \[f(u+v)=f(u)+f(v), f(kv)=kf(v)\] \[\forall u,v \in V,k \in K\] where K is the field, underlying both vector space V, and W. Note that, when you specify a vector space, you have to specify what is the field underlying to that vector space elements k belonging to K, are called scalars

OpenStudy (conqueror):

What is that supposed to mean.... @Michele_Laino

OpenStudy (conqueror):

@iGreen What about linear models...

OpenStudy (michele_laino):

sorry I have made a little error, I rewrite: \[u \in V, v \in W, k \in K\]

OpenStudy (conqueror):

@Michele_Laino Is that supposed to be 9th grade....Because I'm talkin about Algebra 1 9th grade.

OpenStudy (michele_laino):

linear model are used in order to mathematize some physicsl system present in our life: for example a little mass attached to a spring oscillating into a viscous medium. In order to solve that system, is frequently used the Fourier transform. Fourier transform allow us to calculate the response of the system which under examination by the researcher!

OpenStudy (conqueror):

That is still not 9th grade...

OpenStudy (michele_laino):

@Conqueror please tell me something the 9th grade

OpenStudy (conqueror):

like solving this: A semiconductor manufacturing company employs 1,000 people. The average weekly salary of each employee is $750. Select the graph that correctly shows the total cost to the company of paying its employees during a year.